# Probing the Vulnerability of Large Language Models to Polysemantic Interventions

### v1

@@ Abstract
Polysemanticity—where individual neurons encode multiple unrelated features—is a well-known characteristic of large neural networks and remains a central challenge in the interpretability of language models. At the same time, its implications for model safety are also poorly understood. Leveraging recent advances in sparse autoencoders, we investigate the polysemantic structure of two small models (Pythia-70M and GPT-2-Small) and evaluate their vulnerability to targeted, covert interventions at the prompt, feature, token, and neuron levels. Our analysis reveals a consistent polysemantic topology shared across both models. Strikingly, we demonstrate that this structure can be exploited to mount effective interventions on two larger, black-box instruction-tuned models (LLaMA3.1-8B-Instruct and Gemma-2-9B-Instruct). These findings suggest not only the generalizability of the interventions but also point to a stable and transferable polysemantic structure that could potentially persist across architectures and training regimes. Code and data are available here.
@@ 1 Introduction
Polysemanticity refers to the phenomenon in which individual neurons or groups of neurons in neural networks often encode a greater number of distinct features or concepts than the number of neurons involved. This property becomes increasingly prevalent as models scale and has been shown to enhance learning performance  (Wang et al., 2024; Marshall & Kirchner, 2024; Oikarinen & Weng, 2024b). Anthropic’s work on superposition builds on prior insights, showing that large transformer models encode more features than neurons by using linear combinations of activations. This mechanism sacrifices monosemanticity but significantly improves model intelligence  (Elhage et al., 2022). Mathematical analyses reveal that polysemantic neurons enable networks to represent exponentially more features compared to monosemantic approaches  (Elhage et al., 2022).
@@
However, this representational efficiency comes with notable trade-offs. Most significantly, it complicates model interpretability, as entangled representations obscure how human-understandable concepts are encoded within the model’s internal structure. One mechanistic approach to address this challenge is the use of sparse autoencoders (SAEs), which aim to disentangle superimposed features by learning sparse, higher-dimensional representations of model activations. This enables the extraction of more interpretable, monosemantic features, where each SAE neuron ideally corresponds to a single concept (Bricken et al., 2023; Templeton et al., 2024) [^1] Nevertheless, several studies have also documented important limitations of SAEs (see Appendix J).. Recent work has shown that SAE-derived features exhibit a degree of universality across different LLMs (Lan et al., 2024), suggesting the existence of fundamental patterns in how neural networks encode meaning. This consistency hints at the emergence of shared semantic topologies that persist across architectures and training regimes, raising profound questions about whether these patterns are merely computational artifacts or reflections of deeper structures in the world of meanings (Huh et al., 2024). Beyond SAEs, a broader range of interpretability techniques is emerging (Chang et al., 2025; Dunefsky et al., 2024)
@@
The second trade-off, which is largely overlooked in current literature, involves systematic risks stemming from polysemantic structures in language models. In Anthropic’s toy experiments, they briefly note that stronger superposition can make models more vulnerable to adversarial attacks (Elhage et al., 2022). Beyond this, to our knowledge, there is very little existing empirical research that directly addresses the safety implications of polysemanticity in LLMs. In contrast, the vision model domain has a well-established body of work on various forms of attacks that exploit polysemantic representations (Goh et al., 2021; Oikarinen & Weng, 2024a; Geirhos et al., 2023; Dreyer et al., 2024; Huang et al., 2022). Bereska and Gavves, in their review of mechanistic interpretability for AI safety, highlight polysemanticity as a key challenge in building safer LLMs  (Bereska & Gavves, 2024). To address this gap, we focus on polysemantic structures in LLMs—particularly those that persist across models—and explore targeted interventions to better understand their associated risks.
@@
Human Symbolic Manifold ($\mathcal{M}$): The real but largely unobservable space of meanings.
Model Activation Space ($\mathcal{A}_{\ell}$): The $d$-dimensional vector space spanned by the neurons in layer $\ell$ of the LLM; it is the model’s learned approximation of $\mathcal{M}$.
@@
As illustrated in Figure 1, orthogonality in the activation space $\mathcal{A}_{\ell}$ does not persist after projection into the symbolic manifold $\mathcal{M}$. Consequently, two features from $\mathcal{F}_{\ell}$ that appear uncorrelated in $\mathcal{M}$ (i.e., semantically unrelated under human interpretation) can still interfere substantially in $\mathcal{A}_{\ell}$. This interference is also often unevenly distributed across neurons. Building on these two structural vulnerabilities, we design feature, token, prompt, and neuron levels of intervention to investigate: (1) whether model’s expression on a target is sensitive to features and tokens that are semantically unrelated but interfering, and (2) whether model sensitivity correlates with neuron polysemanticity, defined as the number of distinct features a neuron encodes. In this work, model sensitivity to interventions is measured by the shift in the next-token prediction distribution following the intervention.
Our findings are three-fold. Most importantly, we present experimental evidence that interventions leveraging polysemantic structures of LLMs can effectively and covertly alter model outputs. Specifically, by targeting features and tokens—via steering vector techniques—and prompts—via prompt injection—that are not semantically aligned with the intended target but interfere with it, we can reliably induce the model to express the desired semantics. Secondly, we identify the existence of cross-model persistent polysemantic structures. By collecting shared interference features from both Pythia-70M and GPT-2-Small and applying them to steer LLaMA-3.1-8B-Instruct and Gemma-2-9B-Instruct, we still observe substantial intervention effectiveness, revealing a consistent architecture of meaning that transcends specific implementations. Finally, we analyze intervention at the neuron level and find that highly polysemantic neurons are more vulnerable: modifying their activation leads to greater semantic shifts in model output. However, for “super-neurons” (i.e., activated by over $500$ features) amplification strongly alters model behavior, while deactivation has a notably reduced effect, suggesting they may serve as critical junctions in the semantic architecture.
@@ 2.1 Sparse Feature Extraction with SAEs
Our initial exploration of polysemantic structures draws on the pre-trained SAEs provided by Neuronpedia [^2] https://www.neuronpedia.org/. We focus on GPT-2-Small and Pythia-70M, the two models for which Neuronpedia supplies SAEs for the most important sub-modules in every layer. The dimensionality of all the provided SAEs is $32,768$, under which explicit features are extracted. For clarity in subsequent sections, the direction of a feature is defined as its projection into $\mathcal{A}_{\ell}$. The interference between two SAE features in $\mathcal{F}_{\ell}$ is quantified by the cosine similarity of their directions; the semantic relatedness of them is measured by the cosine similarity of their projection into $\mathcal{M}$ instead.
@@ 2.2 Distinct Feature Identification with Agglomerative Clustering
SAEs disentangle polysemantic neurons into monosemantic sparse features, these features, however, are not always decomposed at a consistent semantic level (Bricken et al., 2023; Foote, 2024). For example, a neuron associated with dog-related concepts might be divided into features representing different dog breeds, while another neuron encoding both cat and car concepts might be split into features representing cat and car. In such cases, the resulting features differ in granularity. To mitigate this inconsistency, we employ agglomerative clustering to align feature representations to a consistent semantic level, facilitating both (1) the quantification of neuron polysemanticity and (2) the extraction of semantically distinct feature groups for subsequent analysis.
To achieve distinct, higher-level features, we first compute the semantic relatedness between SAE features, based on their auto-interpretation generated using GPT-4o-mini (Caden Juang et al., 2024). We perform agglomerative clustering in each layer with a relatedness threshold of $0.5$. Figure 2 shows an example of the clustering results for the 5th MLP layer of Pythia-70M. Filtering the connections with strength below $0.2$ and counting the number of aggregated features with which the neurons are aligned, we find that fewer than $5$% of the neurons exhibit polysemanticity in each layer on average, as shown in Figure 15.
Figure 2: Agglomerative clustering of SAE features trained on Pythia-70M’s 5th MLP layer. Only the five largest feature clusters are labeled, and the 10 largest clusters are color-coded.
2.3 Dataset Construction
We construct tailored contextual prompts for each vocabulary token. Specifically, for each token in Pythia-70M and GPT-2-Small, we use DeepSeek-V3 to generate three incomplete sentences with varied contexts in which the token is likely to appear next, resulting in a sentence-completion dataset containing $150,000$ prompts for each model. More concrete considerations and details of our dataset generation are stated in Appendix D. In a word, we conduct interventions in specific contexts to influence the output probability of the target token associated with a SAE feature.
@@ 2.4 Evaluation Criteria
The effectiveness of intervention is evaluated based on the changes of relatedness between the model’s next-token prediction and SAE feature’s top-activating token. In our work, we apply four metrics to quantify the shift in relatedness.
Weighted cosine similarity(c): Similarity $s_{t}$ between a single token $t\in T$ and a token set $\bar{T}$ is computed as the maximum cosine similarity between token embeddings of token $t$ and $\bar{t}\in\bar{T}$. Providing a weight $w_{t}$ for each $t$, the weighted cosine similarity $c(T,\bar{T})$ between two token set $T$ and $\bar{T}$ is then computed as $\Sigma_{t\in T}s_{t}\cdot w_{t}$.
Kendall’s Tau ($\tau$): Kendall’s Tau $\tau(T,\bar{T})$ of two token sets is computed based on the intersection $T\cap\bar{T}$, where token ranks are determined by their positions in $T$ and $\bar{T}$. The absolute value is reported, i.e., $|\tau|$.
Spearman Correlation ($\rho$): Using the same intersection set $T\cap\bar{T}$, we compute Spearman rank correlation coefficient between the ranks of tokens in $T$ and $\bar{T}$. The absolute value is reported: $|\rho|$.
Weighted overlap(w): Providing a weight $w_{t}$ for each $t\in T$, the weighted overlap $w(T,\bar{T})$ of two token sets $T$ and $\bar{T}$ is computed as $\Sigma_{t\in T\cap\bar{T}}w_{t}$.
We limit the scope of our analysis on model’s next-token output to the top-10 predictions, among which the next token is mainly determined. The outputs of model before and after the intervention are then denoted as $O$ and $\tilde{O}$, while the prediction probabilities of tokens are computed as the weight values mentioned above. The token set $T_{t}$ of a target SAE feature to be intervened consisting of the vocabulary tokens with highest cosine similarities of token embeddings with its top-activating token. The effect of intervention thus can be quantified by the improvement in these four metrics before and after conducting the intervention, e.g. $c(\tilde{O},T_{t})-c(O,T_{t})$.
@@ 2.5 Overview of Intervention Methods
Our study of polysemantic intervention starts from Pythia-70M and GPT-2-Small in three different ways: steering with feature direction, steering with token gradient direction, and prompt injection. We first randomly select target features to be intervened [^3] We randomly sample $450$ target features from GPT-2-Small and $1,200$ from Pythia-70M.. For each selected target feature, we sample interference features from feature clusters derived from Section 2.2—excluding the target’s own cluster—to ensure large semantic distance. Interference features are drawn from several predefined interference intervals: $[0.0,0.1]$, $[0.1,0.2]$, $[0.2,0.3]$, $[0.3,0.4]$, and $[0.4,1.0]$. In the previous two studies, the steering vectors are obtained through two ways respectively: employing their feature directions by projecting $\mathcal{F}_{\ell}$ to $\mathcal{A}_{\ell}$ or directions consisting of the gradients computed from the related layer’s neurons with respect to their top-activating tokens. We scale these vectors within the range $[-20,20]$ to maximize improvements in the four metrics stated above. For prompt injection, we inject sampled top-activating tokens of interference features into the prompt and compare the success rate of elevating target features’ top-activating ones to top-10 predicted tokens list across different interference levels. Subsequently, we apply the latter two scalable ways to larger models LLaMA3.1-8B-Instruct and Gemma-2-9B-Instruct, which don’t have pre-trained SAEs. For the target features, we utilize interference features shared between Pythia-70M and GPT-2-Small for intervention. Last but not least, we analyze the impact of neuron polysemanticity on model’s output in Pythia-70M and GPT-2-Small. The strongly connected polysemantic neurons are identified by filtering out connections with strength below $0.2$. Then, the degree of neuron polysemanticity is defined by the number of aggregated features the focal neuron connects to. By suppressing or boosting the activation of neurons with different degrees of polysemanticity, we evaluate the shift of output to the semantics of their aligned aggregated features.
@@ 3.1 Exploiting SAE Feature Directions for Intervention
As described above, we sample interference features for each target feature from different interference levels. More experimental details are elaborated in Appendix F. For each interference-target feature pair, we scale the steering vector to maximize the effectiveness of the intervention. Figure 3 reports the averaged results, and Table 1 shows particular examples.
Figure 3: Effects of activation engineering based on interference feature direction. The x-axis shows steering on the target feature and unrelated features with increasing interference. Error bars show standard errors. ***, **, and * denote t-test significance at $p<0.001$, $p<0.01$, and $p<0.05$ vs. the random baseline (i.e., random feature directions).
For comparison, we include the target feature’s own direction as a steering vector (i.e., with an interference value of 1.0). The results support our hypothesis: steering with semantically unrelated but interfering features can influence the output probability of the target’s top-activating tokens, with stronger effects observed for higher interference values. We also find that SAE-based interventions are generally less effective on GPT-2-Small than on Pythia-70M, likely due to the former’s greater depth, which may dilute the impact of activation changes (Fort, 2023).
Finding 1: Steering with directions of distinct features that interfere with a target can significantly amplify the model’s output on the target feature.

### v2

# Signal in the Noise: Polysemantic Interference Transfers and Predicts Cross-Model Influence

@@ Abstract
Polysemanticity is pervasive in language models and remains a major challenge for interpretation and model behavioral control. Leveraging sparse autoencoders (SAEs), we map the polysemantic topology of two small models (Pythia-70M and GPT-2-Small) to identify SAE feature pairs that are semantically unrelated yet exhibit interference within models. We intervene at four loci (prompt, token, feature, neuron) and measure induced shifts in the next-token prediction distribution, uncovering polysemantic structures that expose a systematic vulnerability in these models. Critically, interventions distilled from counterintuitive interference patterns shared by two small models transfer reliably to larger instruction-tuned models (Llama-3.1-8B/70B-Instruct and Gemma-2-9B-Instruct), yielding predictable behavioral shifts without access to model internals. These findings challenge the view that polysemanticity is purely stochastic, demonstrating instead that interference structures generalize across scale and family. Such generalization suggests a convergent, higher-order organization of internal representations, which is only weakly aligned with intuition and structured by latent regularities, offering new possibilities for both black-box control and theoretical insight into human and artificial cognition. Code and data are available here.
@@ 1 Introduction
Polysemanticity refers to the phenomenon in which individual neurons or groups of neurons in neural networks often encode a greater number of distinct features or concepts than the number of neurons involved. This property becomes increasingly prevalent as models scale and has been shown to enhance learning performance (Wang et al., 2024; Marshall & Kirchner, 2024; Oikarinen & Weng, 2024b). Anthropic’s work on superposition builds on prior insights, showing that large transformer models encode more features than neurons by using linear combinations of activations. This mechanism sacrifices monosemanticity but significantly improves model capability (Elhage et al., 2022). Mathematical analyses reveal that polysemantic neurons enable networks to represent exponentially more features compared to monosemantic approaches (Elhage et al., 2022).
@@
Nevertheless, this representational efficiency comes with trade-offs. Most significantly, it complicates model interpretability, as entangled representations obscure how human-understandable concepts are encoded within the model’s internal structure. One mechanistic approach to address this challenge is the use of sparse autoencoders (SAEs), which aim to disentangle superimposed features by learning sparse, higher-dimensional representations of model activations. SAEs enable the extraction of interpretable, monosemantic features, where each SAE neuron ideally corresponds to a single concept (Bricken et al., 2023; Templeton et al., 2024) [^1] Nevertheless, several studies have also documented limitations of SAEs (see Appendix O).. Recent work has shown that SAE-derived features exhibit a degree of universality across different LLMs (Lan et al., 2024), suggesting the existence of fundamental patterns in how neural networks encode meaning. This consistency hints at the emergence of shared semantic topologies that persist across architectures and training regimes, raising profound questions about whether these patterns are merely computational artifacts or reflections of latent semantic regularities (Huh et al., 2024). Except for SAEs, a broader range of interpretability techniques is emerging simultaneously (Chang et al., 2025; Dunefsky et al., 2024).
@@
The second trade-off, which is largely overlooked in current literature, involves systematic vulnerability stemming from polysemantic structures in language models. In Anthropic’s toy experiments, they note that stronger superposition can make models more vulnerable to adversarial attacks (Elhage et al., 2022). Beyond this, to our knowledge, there is very little existing empirical research that directly addresses the safety implications of polysemanticity in language models. In contrast, the vision model domain has a well-established body of work on various forms of adversarial model control that exploit polysemantic representations (Goh et al., 2021; Oikarinen & Weng, 2024a; Geirhos et al., 2023; Dreyer et al., 2024; Huang et al., 2022). Bereska and Gavves, in their review of mechanistic interpretability for AI safety, highlight polysemanticity as a key challenge in building safer LLMs (Bereska & Gavves, 2024). To bridge this gap, we focus on polysemantic structures in real-world LLMs, particularly those that persist across models, and explore hard-to-detect and targeted interventions to better understand their associated risks.
@@
Human Symbolic Manifold ($\mathcal{M}$): The latent first-order symbolic domain that encodes human-intuitive semantics independent of contextual usage.
Model Activation Space ($\mathcal{A}_{\ell}$): The $d$-dimensional vector space spanned by the neurons in layer $\ell$ of the language model; it partially reflects $\mathcal{M}$.
@@
As illustrated in Figure 1, orthogonality in the activation space $\mathcal{A}_{\ell}$ does not persist after projection into the symbolic manifold $\mathcal{M}$. Consequently, two features from $\mathcal{F}_{\ell}$ that appear unrelated in $\mathcal{M}$ (i.e., anchoring semantically to distinct meanings under interpretation) can still interfere substantially in $\mathcal{A}_{\ell}$. This interference is also often unevenly distributed across neurons. Building on these two structural characteristics, we design feature, token, prompt, and neuron levels of intervention to investigate: (1) how models’ expression on a target feature is affected if semantically unrelated but interfering features are manipulated, (2) whether model vulnerability correlates with neuron polysemanticity, defined as the number of distinct features a neuron encodes, and (3) to what extent these polysemantic interference patterns transfer across models. In this work, model vulnerability to interventions is measured by the shift in the next-token prediction distribution following the intervention.
Our findings are four-fold. First, we present experimental evidence that interventions leveraging polysemantic structures of LLMs can effectively manipulate model outputs. Specifically, by targeting features and tokens — via steering vector techniques — and prompts — via prompt injection — that are not semantically aligned with the intended target but interfere with it, we can reliably induce the model to express the desired semantics. Second, we identify the existence of cross-model persistent polysemantic structures. By collecting shared interference features from both Pythia-70M and GPT-2-Small and applying them to steer Llama-3.1-8B/70B-Instruct and Gemma-2-9B-Instruct, we still observe substantial intervention effectiveness, revealing a consistent architecture of meaning that transcends specific implementations. Third, we explore those counterintuitive yet stable interference patterns that replicate across models. Post-hoc annotation of higher-order semantic relations accounts for a minority of cases, indicating that models learn robust regularities largely opaque to interpretation. Fourth, we analyze intervention at the neuron level and find that highly polysemantic neurons are more vulnerable: modifying their activation leads to greater semantic shifts in model output. For “super-neurons” (i.e., activated by over $500$ features), amplification strongly alters model behavior, while deactivation has a notably reduced effect, suggesting they may serve as critical junctions in the semantic architecture.
@@ 2.1 Sparse Feature Extraction with SAEs
Our initial exploration of polysemantic structures draws on the pre-trained SAEs provided by Neuronpedia [^2] https://www.neuronpedia.org/. We focus on GPT-2-Small and Pythia-70M, the two models for which Neuronpedia supplies SAEs for all important sub-modules in every layer. The dimensionality of all the provided SAEs is $32,768$, under which explicit features are extracted. For clarity, in subsequent sections, the direction of an SAE feature is defined as its projection into $\mathcal{A}_{\ell}$. The interference scale between two SAE features is quantified by the cosine similarity of their projected directions in $\mathcal{A}_{\ell}$; their surface-level semantic resemblance is measured by the cosine similarity of their projections into $\mathcal{M}$.
@@ 2.2 Distinct Feature Identification with Agglomerative Clustering
SAEs disentangle polysemantic neurons into monosemantic sparse features. These features, however, are not always decomposed at a consistent semantic level (Bricken et al., 2023; Foote, 2024). For example, a neuron associated with dog-related concepts might be divided into features representing different dog breeds, while another neuron encoding both cat and car concepts might be split into features representing cat and car. In such cases, the resulting monosemantic features differ in granularity. To mitigate this inconsistency, we employ agglomerative clustering to align feature representations to a consistent semantic level, facilitating both (1) the quantification of neuron polysemanticity and (2) the isolation of feature groups exhibiting low similarity in their surface meanings for subsequent analysis.
To identify distinct, higher-level features, we compute the semantic similarity between pairs of SAE features using their auto-interpretation glosses generated by GPT-4o-mini (Caden Juang et al., 2024) [^3] Because different models can yield divergent auto-interpretations of the same feature, we conduct a cross-validation with DeepSeek-V3 on a single SAE layer of Pythia-70M, which reveals both strong concordance between the interpretations produced by GPT-4o-mini and DeepSeek-V3. Replicating experiments with DeepSeek-V3 feature glosses further confirm the consistency of our principal findings. See Appendix J. together with embeddings of the feature glosses from text-embedding-3-large. Prior work offers different heuristics for identifying semantically distinct SAE features. Foote et al. propose a cutoff of $0.5$ for distinguishing semantically distinct feature clusters (Foote, 2024), while another work on analyzing text-embedding-3-large shows that unrelated concepts typically fall within the 0.050.05-$0.30$ cosine similarity range (Zuchen et al., 2025). Drawing on these insights, we conduct agglomerative clustering in each SAE layer using four increasingly strict similarity cutoffs (i.e., $0.40$, $0.30$, $0.20$, and $0.15$) to assess whether our results are robust across varying thresholds of semantic distinctness, while retaining sufficient feature density for experimentation. Figure 2 shows an example of the clustering results for the 5th MLP layer of Pythia-70M under the similarity cutoff of $0.4$. Detailed descriptive statistics on (1) the distribution of cosine similarities among SAE features, (2) the distribution of their interference values, and (3) the correlation between interference scale and semantic similarity are provided in Appendix F.
Figure 2: Agglomerative clustering of SAE features trained on Pythia-70M’s 5th MLP layer under the cosine similarity threshold of $\mathbf{0.4}$. Only the five largest feature clusters are labeled.
@@ 2.3 Evaluation Criteria
The effect of a polysemantic intervention is quantified as the change in alignment between the model’s next-token prediction distribution and a target SAE feature $f\in\mathcal{F}_{\ell}$ on a context sentence. Details about sentence generation are elaborated in Appendix E. Specifically, we assess the similarity between the model’s output and a feature-associated token set $T_{f}\subset V$, which consists of tokens with activation values above a threshold (here we use 0.8 $\cdot$ highest activation value).
Let $O,\tilde{O}\in\Delta^{|V|}$ be the model’s output distributions before and after intervention, and let $E\in\mathbb{R}^{|V|\times d}$ denote the token embedding matrix. Our main metric is weighted cosine similarity:
$c(P,T_{f})=\sum_{t\in V}P(t)\cdot\max_{\bar{t}\in T_{f}}\cos(E_{t},E_{\bar{t}}),$ (1)
where each token $t\in V$ contributes its predicted probability weighted by the highest cosine similarity between its embedding and those in $T_{f}$. This captures how semantically aligned the model’s output distribution is with the target feature direction. Then, the intervention effect is:
$\Delta c=\frac{c(\tilde{O},T_{f})-c(O,T_{f})}{c(O,T_{f})}.$ (2)
As an alternative, we also report weighted overlap, which directly sums the output probability mass over $T_{f}$. Formal definition and results are provided in Appendix G.
@@ 2.4 Overview of Intervention Methods
Our investigation of polysemantic interventions begins with Pythia-70M and GPT-2-Small, using three complementary approaches: feature-direction steering, token-gradient steering, and prompt injection. We randomly select target features to be intervened [^4] We randomly sample $480$ target features from GPT-2-Small and $180$ from Pythia-70M (See Appendix H for details).. For each selected target feature, we sample interference features from feature clusters derived from Section 2.2—excluding the target’s own cluster—to ensure sufficient meaning dissimilarity with the target. Interference features are drawn from five interference intervals: $[0.0,0.1]$, $[0.1,0.2]$, $[0.2,0.3]$, $[0.3,0.4]$, and $[0.4,1.0]$. In the first two experiments, we construct steering vectors for interference features using two methods: (1) projecting feature directions from the sparse basis $\mathcal{F}_{\ell}$ into the model’s activation space $\mathcal{A}_{\ell}$, and (2) computing token-gradient directions from the partial derivatives of the layer’s activations with respect to each feature’s top-activating tokens. For each intervention, we roughly optimize the scaling of the steering vector over the range [20,20][-20,20], balancing intervention strength with the need to preserve coherent model outputs (i.e., avoiding substantial disruption to the overall output distribution). Details about the tuning strategy are elaborated in the Appendix I.2.
In the prompt injection setting, we prepend sampled top-activating tokens from interference features to the input and, rather than using weighted cosine similarity, measure the frequency with which the tokens of target type appear among the model’s top predictions, conditioned on varying levels of feature interference. To evaluate cross-model transferability, we apply the two scalable intervention methods (i.e., token-gradient steering and prompt-injection) to black-box, larger models Llama-3.1-8B/70B-Instruct and Gemma-2-9B-Instruct. For these interventions, we use as targets the interference features shared between Pythia-70M and GPT-2-Small. One might question the practical significance of our next-token control test. While the primary aim of this study is mechanistic, we also include a small-scale evaluation of gradient-based intervention on the HellaSwag dataset to assess its ability to steer model behavior in specific target-related cases without compromising general performance. Details are explained in Appendix L.
Finally, we analyze the impact of neuron polysemanticity on model outputs in Pythia-70M and GPT-2-Small. For each neuron, we identify strongly connected features by thresholding connection weights at $0.2$, and define its degree of polysemanticity as the number of such features. We then suppress or amplify the activations of neurons with varying polysemanticity levels and evaluate how the model’s output distribution shifts toward the semantics of associated feature clusters.
@@ 3.1 Exploiting SAE Feature Directions for Intervention
Figure 3: Effects of intervention based on the interference SAE feature direction. (A–B) Relative change in weighted cosine similarity toward the target ($\Delta c$). Bars show the mean relative change compared to baseline across interference levels, with lighter shades indicating stricter feature-meaning relevancy cutoff thresholds. The x-axis denotes the interference scale between the target and intervention feature: Original corresponds to intervening with the target feature itself, and Random serves as a random feature intervention baseline. Error bars denote 95% confidence intervals. (C) Regression estimates of the effect of feature-pair interference value on intervention success. Two regression specifications are shown: Model A regresses weighted cosine similarity after intervention ($c(\tilde{O},T_{f})$) on interference value, with feature-meaning similarity, baseline weighted cosine similarity ($c(O,T_{f})$), and layer-type controls; Model B regresses the change score ($\Delta c$) on interference value, with feature-meaning similarity and layer-type controls. Error bars denote 90%, 95%, and 99% confidence intervals. Results with the alternative metric are shown in Figure 9.
Our hypothesis posits that if two feature directions interfere in $\mathcal{A}_{\ell}$, despite being nearly orthogonal in $\mathcal{M}$, then enhancing one will inevitably influence the other to some degree. If true, this would imply the potential to covertly manipulate the output probability of a target feature by steering with not obviously related features. To evaluate this, we pair each target feature with interference features sampled at varying levels of semantic dissimilarity. We vary feature irrelevance thresholds (from $0.4$ to $0.15$; Appendix H) to confirm that the observed pattern holds across threshold choices. Additionally, we perform controlled regressions—accounting for feature-pair cosine similarity—under two linear model specifications to confirm that interference effects persist independently of semantic similarity. Figure 3 reports the results, and Table 3 in Appendix H shows particular examples.
All analyses consistently show that steering with features that are semantically dissimilar yet interfering can alter the output probabilities of a target feature’s top-activating tokens, with stronger effects observed at higher interference levels. We also find that SAE-based interventions are generally much less effective on GPT-2-Small than on Pythia-70M, likely due to the greater depth of the former, which may attenuate the influence of mid-layer activation shifts (Fort, 2023).